How do you evaluate the six trigonometric functions of theta.
a. theta= 90
b. theta= -1pi

Answer by stanbon(75887)

How do you evaluate the six trigonometric functions of theta.
a. theta= 90
------
Look at a unit circle centered on (0,0).
90 degrees intersects the circle at (0,1)
x = 0 ; y = 1, and r = 1 because you have a unit circle
sin(t) = y/r = y = 1
cos(t) = x/r = x = 0
tan(t) = y/x = 1/0 is undefined
-----------------
csc, sec, and cot are the inverse of sin, cos, tan
----------------------
b. theta= -1pi
The point on the unit circle is (-1,0)
x = -1, y = 0, r = 1
sin(t) = y = 0
cos(t) = x = -1
tan(t) = y/x = 0
---
csc, sec, and cot are the inverse of sin, cos, tan
=======================
cheers,
Stan H.